OpenStudy (anonymous):

Solve the equation: 2x+10=30 :)

OpenStudy (anonymous):

1.subtract 10 on both sides 2.divide 2 on both sides

OpenStudy (sasogeek):

x=10

OpenStudy (sasogeek):

2x=30-10 x=20/2 x=10

OpenStudy (anonymous):

Thanks!!